### v1

# Compensation behaviour in trilayered anisotropic 6-state clock model

Abstract
The equilibrium behaviours of a trilayered 6-state clock model has been investigated by Monte Carlo simulation. The intralayer interaction is considered ferromagnetic, whereas the interlayer interaction is antiferromagnetic. Open boundary conditions are applied in all directions. The thermodynamic behaviours of sublattice magnetisation, total magnetisation, magnetic susceptibility and the specific heat are studied as functions of the temperature. The interesting compensation phenomenon, i.e., the vanishing of the total magnetisation even with nonvanishing sublattice magnetisation, has been observed. The compensation temperature and the critical temperature are found to depend on the strength of single-site anisotropy. The comprehensive phase diagram is shown in the temperature-anisotropy plane. The finite size analysis has been done through the temperature dependence of fourth order Binder cumulant, and the scaling exponents are estimated.
Keywords: Clock model, Anisotropy, Monte Carlo methods, Sublattice magnetisation, Compensation, Binder cumulant
PACS Nos: 05.50.+q, 05.10.Gg, 64.60.-i, 75.30.Gw, 75.30.Ds, 75.10.-b,
E-mail:mallickolivia0@gmail.com
+E-mail:muktish.physics@presiuniv.ac.in
1 Introduction
The ferromagnetic models showing the unusual ordered phase even in absence of conventional long-range ferromagnetic order is a remarkable discovery[1, 2]. The continuous symmetric (SO(2)) planar ferromagnetic model shows this behaviour in two dimensions with vortex-antivortex special kind of ordered state[3, 4, 5]. The unconventional metastable behaviour and phase transition in this kind of model has drawn the attention of the researchers in last four decades. The celebrated BKT transition is now under the spotlight of modern research of condensed matter physics and statistical mechanics.
Is this kind of unusual phase the property of continuous symmetry only? This question has triggered to investigate the discrete symetric q-state clock model, where the continuous symetric XY model can be restored in the limit $q\to\infty$. The properties of q-state clock model has been investigated[6] for q=4, 5 and 6 in two dimensions by Monte Carlo renormalization group analysis to study the phase transitions. The Kosterlitz-Thouless type intermediate phase has been found[7] in two dimensional 6-state clock model by Monte Carlo simulations. The KT type phase transition has been found to occur in the range between ${{k_{B}T}\over{J}}=0.68\pm 0.02$ and ${{k_{B}T}\over{J}}=0.92\pm 0.01$. However, it is confirmed later that there exists a Kosterlitz-Thouless (KT) like phase in the ferromagnetic 6-clock model on the square lattice through studies of the interfacial free energy estimated by Monte Carlo simulations and authors estimated[8] that lower and upper transition temperatures are 0.75 and 0.90 respectively. Interestingly, in a recent study[9], the machine-learning detection of BKT transition was found in q-state clock model with the neural network based estimates of the two critical points as ${{k_{B}T}\over{J}}=0.410$ and ${{k_{B}T}\over{J}}=0.921$.
The classical 2D clock model is known to have a critical phase with Berezinskii–Kosterlitz–Thouless (BKT) transitions. These transitions have logarithmic corrections which make numerical analysis difficult. Hence, the classical 2D six-state clock model has been studied[10] by mapping it to the quantum self-dual 1D six-state clock model and confirmed that the self-dual point has a precise numerical agreement with the analytical result. It has been argued also that the degeneracy of the excited states at the self-dual point from the effective field theoretical point of view.
The thermodynamic behaviours of the clock model has been studied in digfferent kinds of lattices to investigate the effects of geometrical connections and interactions. The BKT transition has been observed[11] in 6-state clock model on the rewired squared lattice studied by Monte Carlo simulation with Wang-Landau algorithm and estimated the lower and upper critical temperatures. For pure case, i.e., the lattice with average number of bonds equal to 2, ${k_{B}T1}\over{J}$} and ${k_{B}T2}\over{J}$}, are roughly estimated around 0.55 and 1.15 respectively (these results may be compared to those reported in the references [7] [8] and [9]). The BKT transition has been studied[12] recently, in 6-state clock model on different kinds (triangula, kagome, honyecomb and diced) of two dimensional lattices. The two dimensional q-state clock model on random lattices has been investigated[13] via numerical simulation. Interestingly, for q=4 a single phase transition has been reported. Whereas, for q=5 and q=6 multiple (double) phase transitions are observed from the two peaks of the specific heat. The lower and upper critical temperatures are found to depend significantly on the structure of the lattice. The anisotropic (with anisotropy parameter $\alpha$) 6-state clock model has been investigated[14] by transfer matrix renormalization group method to explore the phase transition in two dimensional square lattice. Three distinct phase are identified and a comprehensive phase diagram is drawn. They observed a single discontinuous phase transition and two BKT phase transitions meeting in the tricritical point. The tricritical point is estimated at critical anisotropy $\alpha_{c}=0.21405$. The antiferromagnetic 5-state clock-model has been studied on the triangular lattice recently[15] with second nearest neighbour interaction and the phase diagram is drawn.
The thermodynamic behaviours of clock model in layered structured lattices are not yet reported in the well developed literature of this specific model. However, the behaviours of layered Ising and Blume-Capel models are well studied. Let us briefly mention a few selected studies on layered Ising and Blume-Capel models. The main interesting phenomenon in the thermodynamic behaviours of layered Ising and Blume-Capels models is the compensation transition. The compensation temperatures and magnetic susceptibility of a mixed ferro-ferrimagnetic ternary alloy has been studied[16] by Monte Carlo simulations and observed outstanding behaviour of magnetic susceptibility near the compensation and critical temperature. The compensation temperature has been found[17] to depend of the anisotropy in a mixed spin (spin-${1}\over{2}$} and spin-$1$) Heisenberg ferrimagnet on a simple cubic lattice, studied by Oguchi approximation. The anisotopy dependent spin compensation temperature has also been noticed[18] in mixed spin (spin-${3}\over{2}$} and spin-${5}\over{2}$}) ferrimagnetic Ising model studied by mean field approximation. In this case, the possibility of multiple compensation temperature was indicated. The impurity- concentration dependent compensation[19] and the anisotropy dependent compensation[20] was also noticed in the trilayered Ising and Blume-Capel models, respectively. The meanfield and effective field theory has been employed[21] to study the interaction dependent compensation in spin-${1}\over{2}$} Ising trilayered ferrimagnetic system. The phase diagram (compensation and no-compensation) aws drawn in parameter space. The antiferromagnetic systems showed the compensation behaviours in Ising mixed spin (spin-${3}\over{2}$} and spin-${5}\over{2}$}) core/shell nanowire[22] and in mixed spin (spin-${7}\over{2}$} and spin-1{1}) antiferromagnetic ovalene nanowire[23]. The compensation wad found[24] Ising mixed spin (spin-${3}\over{2}$} and spin-${1}\over{2}$}) graphene bilayered ferrimagnet. The effects of exchange interactions, crystal field and external magnetic field on the compensation temperature was studied[24] by Monte Carlo simulation. The evidence of engineering the compensation in ferrimagnetic iron garnet by tuning the ratio of in-plane and out of plane magnetisations can be found in [25]. An interesting spin compensation and first order reentrant behaviour has been observed[26] in mixed spin (spin-1 and spin-2) Ising ferrimagnetic system studied by Monte Carlo simulation.
Can the discrete symmetric planar models (6-state clock model) show the compensation behaviour? Being motivated by this question, we have thoroughly investigated the trilayered 6-state clock model by Monte Carlo simulation. The results are reported with diagrams in this manuscript. This manuscript is organized as follows: The model and the simulational scheme are discussed in Section-2. The numerical results are reported with diagrams in Section-3 and the paper ends with concluding remarks in Section-4.
2 Model and numerical details
We consider a classical trilayered 6-state clock model defined on a simple cubic lattice, where each of the three monoatomic layers—denoted $l_{1}$, $l_{2}$, $l_{3}$—consists exclusively of either type-A or type-B atoms. The magnetic interactions and anisotropic contributions to the energy of the system are captured by the following Hamiltonian:
$\mathcal{H}=-J_{ferro}^{intra}\sum_{\langle ij\rangle}\vec{S_{i}}\cdot\vec{S_{j}}-J_{antiferro}^{inter}\sum_{\langle ij\rangle}\vec{S_{i}}\cdot\vec{S_{j}}-D\sum_{i}[({S_{i}}^{x})^{2}-({S_{i}}^{y})^{2}]$ (1)
Here, $\vec{S_{i}}=(S_{i}^{x},S_{i}^{y})$ represents the classical two dimensional spin vector at lattice site $i$, which can point in one of six discrete directions in the $xy$-plane, consistent with the 6-state clock model. These discrete orientations are defined by:
$S_{i}^{x}=\cos{\theta_{i}},S_{i}^{y}=\sin{\theta_{i}},$
where $\theta_{i}$ can assume any value discretely,
$\theta_{i}=\frac{2\pi}{6}n_{i},[n_{i}=0,1,2..5]$
The first term in Eq-(1) accounts for the intralayer ferromagnetic interactions between nearest neighbour spins $\langle ij\rangle$ within the same atomic plane, with a positive coupling exchange $J_{ferro}^{intra}>0$. The second term represents interlayer antiferromagnetic coupling between the adjacent layers, involving nearest neighbour sites $\langle ij\rangle$ across the layers, governed by a coupling constant $J_{antiferro}^{inter}<0$. The third term introduces an anisotropy D, which destroys the discrete rotational symmetry in the $xy$-plane and favours the spin-alignment in specific direction. Positive values of $D$ prefer specific spin orientations, thereby reinforcing the discrete six-fold symmetry and enhancing the stability of the selected clock states. We impose periodic boundary conditions (PBC) in the x and y directions within each layer and along the stacking direction z; we adopt free boundary conditions to capture the finite thickness of the trilayered structure.
To study the compensation behaviour in the trilayered 6-state clock model, we employ Monte Carlo simulation technique based on Metropolis algorithm with random single-spin updates. The system consists of three stacked square lattices, each consisting of $L\times L$ sites, resulting in a total of $3L^{2}$ spins. For majority of our simulations, we consider a system size of $L=128$, corresponding to $128\times 128$ spins per layer. In order to examine finite-size effects, additional simulations are performed for various lattice sizes. Initially, spin configurations are generated at high temperature with randomly assigned orientations chosen from the six allowed states in the clock model. Then the system is allowed to evolve using Metropolis updating scheme: during each Monte Carlo steps (MCS), $L^{2}$ spin-flip attempts are made, where a trial spin selected at random and updated according to the Metropolis acceptance probability[27]: $P_{f}=Min[\exp{-\frac{\delta E}{k_{B}T}},1]$. $\Delta E$ is the change in energy associated with the proposed spin flip, and $k_{B}(=1)$ is the Boltzmann constant. The simulation spans a duration $2\times 10^{5}$ MCSS within an initial transient segment of $10^{5}$ MCSS excluded to achieve the assumed ergodic limit. To characterize the thermodynamic behaviour of the 6-state clock model, we calculate several key observables that capture the collective magnetic ordering of the system. The sublattice magnetisation provides insight into the local symmetry breaking and is defined for each sublattice $k\in{1,2,3}$ as the average spin components in the x and y directions:
$m_{k}^{x}=\frac{1}{L^{2}}\sum_{i,j}S^{x}(i,j,k)\hskip 14.22636ptm_{k}^{y}=\frac{1}{L^{2}}\sum_{i,j}S^{y}(i,j,k)$
where $S^{x}(i,j,k)$ and $S^{y}(i,j,k)$ denote the Cartesian components of the spin vector on the site $(i,j)$ belonging to the $k$-th sublattice and $L^{2}$ is the total number of lattice site in each sublattice. The total magnetization of the system is obtained by summing the contributions from all sublattices:
$M=\sqrt{M_{x}^{2}+M_{y}^{2}}$
where,
$M_{x}=\sum_{k}m_{k}^{x}\hskip 14.22636ptM_{y}=\sum_{k}m_{k}^{y}$
We compute corresponding magnetic susceptibility by
$\chi=\frac{3L^{2}}{k_{B}T}(\langle M^{2}\rangle-\langle M\rangle^{2})$
where $T$ is the temperature and $k_{B}$ is Boltzman constant. The specific heat $C$ is defined as ,
$C=\frac{d\langle E\rangle}{dT}$
where $\langle E\rangle$ is the average energy density calculated from the Hamiltonian. We also calculate the thermal variation of fourth-order Binder cumulant for different system sizes,
$U_{L}=1-\frac{\langle M^{4}\rangle}{3\langle M^{2}\rangle^{2}}$
3 Results and Discussion
3.1 Isotropic regime:$D=0$
First, we discuss the thermodynamic phase transition and compensation phenomenon of a trilayered 6-state clock at zero anisotropy($D=0$). All six clock states are energetically favourable in this regime, as there is no directional bias. The intralayer ferromagnetic coupling is taken for A-type atoms as $J_{ferro}^{AA}=0.2$ and for B-type atoms as $J_{ferro}^{BB}=1.0$. The interlayer antiferromagnetic coupling strength is taken as $J_{antiferro}^{AB}=-0.1$. Initially, the system is at high-temperature disordered paramagnetic state, characterised by random spin orientations and zero magnetisation. The system is slowly cooled down with small temperature step $dT=0.05$. As the temperature decreases, magnetic ordering sets in. Figure-1(a) shows the growth of total magnetization $M(T)$, which increases continuously from zero to a non-zero finite value at a well-defined critical temperature $T_{c}$. The transition temperature $T_{c}$ is precisely identified from the position of the peak of the magnetic susceptibility $\chi(T)$ as shown in Figure-1(b).
Interestingly, below $T_{c}$, the total magnetisation again vanishes at a lower temperature $T_{comp}$, referred to as compensation temperature. This occurs even though the sublattices (individual layers) remain partially ordered. Compensation is a peculiar balancing situation where the total magnetization of all layers vanishes, despite the nonzero sublattice ordering. The susceptibility curve $\chi(T)$ exhibits a secondary peak at this temperature, further supporting the presence of a sharp magnetic cross-over. The corresponding specific heat $C(T)$ shown in Figure-1(c), also reflects these thermodynamic features with distinct peaks at both $T_{c}$ and $T_{comp}$. The compensation phenomenon originates from the competing magnetic interations: each layer offers ferromagnetic intra-layer coupling, favouring uniform spin alignment within the layer, where adjacent layers are coupled via anti-ferromagnetic interlayer interactions leading to repel layer magnetisation.
To acquire a more nuanced perspective to the magnetic ordering in each layer, we analyse the sublattice magnetisation of the individual layers across the temperature. Figure-2 shows the sublattice magnetisation components $m_{x}$ and $m_{y}$ as functions of temperature for three distinct layers. These profiles describe the growth of magnetic ordering in each layer. At $D=0$, where no easy-axis anisotropy is present, the system does not energetically favour in any specific direction in the $xy$ plane. Consequently, both $m_{x}$ and $m_{y}$ develop finite values below the transition temperature, indicating spontaneous symmetry breaking. As shown in Figure-2(a) or (b) all three layers exhibit ferromagnetic ordering below the transition temperature $T_{c}$ with each layer developing finite magnetisation. However, due to the antiferromagnetic interlayer coupling, the magnetisations of adjacent layers are oriented in opposite directions. In particular, the middle layer tends to be antiparallel to the outer layers, leading to cancellation of total magnetisation at $T_{comp}$.
To visualise the spin ordering, we also examine the lattice morphology at representative temperatures. Spin configurations snapshots displayed in Figure-3 reveal a transition from a disordered paramagnetic state at high temperature to a well-aligned domain at low temperature for different layers. At very high temperature $T>>T_{c}$ spin are randomly oriented in all three layers. Six colours denote the six specific clock states. For the isotropic regime $(D=0)$, at high temperature all six directions are equally preferable. As temperature decreases,the spins at each layer progressively align, leading to the nucleation and growth of the magnetic domain, which marks the onset of long-range order and results a non-zero total magnetization. Near $T_{comp}$, distinct domain patterns emerge with opposing magnetisation contributions, reflecting the layer-wise compensation behaviour. To quantatively validate the spin morphology, we analysed the angular distribution of spins across the three layers at different temperature. The normalized spin population (fraction of total spins) is plotted as a function of six discrete clock states ( corresponding spin angle 0, $\pi/3$, $2\pi/3$, $3\pi/3$, $4\pi/3$, $5\pi/3$) for each layer. At high temperature, the spin population is nearly uniform across all angles in all three layers, reflecting an isotropic regime where all spin states are equally probable. As temperature decreases towards the critical value, certain spin angles become more populated, indicating the onset of preferred aligment. Near compensation temperature, layer 1 and 3 exihibit a dominant population at $\theta=\pi/3$(n=2), whereas layer 2 prefers $\theta=5\pi/3$. Importantly, the other spin angles are populated such that net magnetization of the whole system becomes zero, while each layer individually retains a finite magnetisation. Upon further cooling to low temperature, layers 1 and 3 show nearly complete alignment at $\theta=\pi/3$ while layer 2 remains oriented in $5\pi/3$. The visual and quantitative results together confirm the layered character of magnetic ordering and the competing interactions that drive the compensation phenomenon.
3.2 Anisotropic regime: $D>0$
In this section, we explore the anisotropic response of the clock spins. By varying the anisotropy parameter D, we systematically investigate the evolution of key observables such as magnetisation, susceptibility and specific heat. Figure-7(a) shows the thermal variation of total magnetisation as function of temperature for different anisotropy values. As temperature decreases, magnetisation grows, exhibiting a clear phase transition at $T_{c}$. Below $T_{c}$, the total magnetisation again vanishes at a lower temperature, denoted as the compensation point $T_{comp}$, indicating the compensation behaviour. The critical temperature $T_{c}$ and compensation temperature $T_{comp}$ are precisely obtained from the peak positions of magnetic susceptibility curves Figure-7(b). The corresponding specific heat profiles further corroborate these findings by exhibiting peaks near both temperatures.
Now, if we increase D, the susceptibility peak shifts toward the right, indicating the critical temperature increases. $T_{comp}$ also shifts slightly. This is captured in the phase diagram in D-T plane shown in Figure-9. This diagram summarizes the evolution of both the thermodynamic phase transition temperature $(T_{c})$ and the compesation temperature $(T_{comp})$ as a function of anisotropy $D$. A monotonic increase in $T_{c}$ observed with increasing $D$, indicationg that stronger single-site anisotropy stabilizes the ordered phase and shifts the thermodynamic transition to higher temperatures. In contrast, $T_{comp}$ exhibit only a slight upword trend with increasing $D$ and remains well separated from $T_{c}$ across the probed range. The critical temperature $T_{c}$ is prescribed by collective critical fluctuations, which are highly sensitive to the anisotropy parameter $D$. Increasing $D$ enhances spin alignment and supressess thermal disorder, thereby stabilizing the ordered phase and shifting $T_{c}$ to higher temperatures. Besides, the compensation temeperature $T_{comp}$ does not correspond to bonafide thermodynamic singularity. Instead it arises from a subtle cancellation between the magnetizations of individual sublattices. This balance dependes primarily on competing intra and inter layer exchange interactions and local spin configurations rather than on long-range critical fluctuations, resulting in only a weak variation of $T_{comp}$ with D.
We also analyze the sublattice magnetization of different layers for different strengths of anisotropy. As positive $D$ influences magnetic ordering along $x$ direction, $y$ component of magnetisation $m_{y}$ remains zero always.As we decrease the temperature $m_{x}$ grows in each of the layers, resulting a thermal phase transition of the total system at a certain temperature $T_{c}$. This is shown in Figure-7. Due to the antiferromagnetic coupling between the adjacent layers, magnetic ordering are opposite in the respective layers. As a result, at a certain temperature ($T_{comp}$) they cancel each other, total magnetization becomes zero. But the individual layers possess a net magnetic moment. The compensation temperature $T_{comp}$ also increases (with small variation) as the strength of anisotropy increases.
To visualize the spin ordering, we also study the spin morphology at different temperature for anisotropy $D=2.0$. This is depicted in Figure-8.At $T>>T_{c}$ the spins are randomly oriented at all layers, resulting in zero magnetisation. Near the critical temperature $T_{c}$ , thermal fluctuations are still strong enough to disrupt complete spin alignment, but a tendency emerges for spins to align along either the positive or negative $x$ direction. This marks the onset of magnetic domain growth, where regions of correlated spins begin to form and compete.
At the compensation point $T_{comp}$, the system exhibits a striking balance: the spins in the two outer layers align in one direction, while those in the middle layer orient in the opposite direction. As a result, the net magnetization of the entire system cancels out to zero, even though each layer retains a finite magnetization. This phenomenon is characteristic of ferrimagnetic-like ordering, where sublattice magnetizations compensate each other.
At very low temperatures, the system reaches a fully ordered state. In this regime, the spins in the two outer layers are aligned along the negative $x$-direction, while the spins in the middle layer are aligned along the positive x-direction. This configuration reflects the combined effects of interlayer exchange interactions and anisotropy. To further elucidate the role of anisotropy in shaping the spin morphology, we also examined the angular distribution of spins for the case of $D=2.0$, where the easy-axis energetically favors spin orientations along $\theta=0$ or $\theta=\pi$. This is shown in Figure- . At high temperature, the majority of spins in all layers are distributed near $\theta=0$, and $\theta=\pi$ reflecting the dominant easy-axis alignment imposed by the positive anisotropy. Upon cooling, the distribution becomes progressively more ordered, with the relative populations $\theta=0$ and $\theta=\pi$ starting to compete. Near the compensation temperature, a distinct redistribution occurs: layers 1 and 3 exhibit a stronger spin population at $\theta=\pi$, whereas layer 2 retains it alignment at $\theta=0$. This antiparallel arrangement of sublattices lead to a vanishing total magnetisation while preserving finite sublattice magnetisation. At very low temperature, the system reaches a nearly complete alignment where layers 1 and 3 are oriented along $\theta=\pi$ and layer 2 aligned along $\theta=0$. At sufficiently low temperatures, the system recovers a finite total magnetisation as thermal fluctuation subside. Thus, while both the isotropic ($D=0$) and anisotropic ($D=2.0$) systems exhibit a compensation behaviour governed by the redistribution of opposing spin populations, the positive anisotropy merely selects preferred axes ($\theta=0,\pi$) without changing the underlying mechanism.
3.3 Finite size analysis
Finite-size effects play a crucial role in accurately characterising the thermodynamic phase transitions in numerical simulations. To investigate the finite-size behaviour of the clock model, we performed simulations for different system sizes $L=16,32,64,128,256$. The fourth-order Binder-cumulant $U_{L}(T)$, magnetisation $M(T)$ and corresponding susceptibilty $\chi(T)$ are studied as a function of temperature. The numerical results are shown in Figure-10 for the considered system sizes. Figure-10(a) illustrates the Binder Cumulant $U_{L}(T)$ where a well-defined intersection provides a robust indicator of the the critical temperature, as it is largely independent of $L$ near the thermodynamic transition. Figure-10(b) shows the magnetization, revealing the expected order-disorder behaviour as the temperature crosses the transition. The susceptibility curves for different system sizes highlights two distinct features: a prominant peak at the thermodyamic critical temperature and a secondary peak at the compensation point. From the Binder cumulant analysis, the intersection point of different $L$ curves yields a critical temperature $T_{c}=1.855$ marking the thermodynamic phase transition at anisotropy $D=2.0$. The susceptibility peak at $T_{c}$ exhibits a systematic increase with increasing system sizes, confirming the growth of critical correlation for the second-order phase transition. The susceptibility diverges in the thermodynamic limit ($L\rightarrow\infty$) according to the scaling $\chi_{max}\sim L^{\frac{\gamma}{\nu}}$. In contrast, the secondary peak associated with the compensation point does not show such scaling behaviour; its height remains nearly constant across the studied sizes. This suggest that the compensation phenomena is not a critical transition, but rather a finite, non-divergent reorganization of the magnetic sublattices.The finite size analysis confirms that the clock model undergoes a thermodynamic phase transition at $T_{c}$ , where critical fluctuations become system-size dependent, while the compensation point reflects a non-critical feature, a thermal anomaly.
4 Concluding remarks
The magnetic compensation behaviour (vanishing of the total magnetisation for nonzero sublattice magnetisation) in a trilayered ferrimagnetic system is an interesting effect studied in discrete symmetric spin models (e.g., Ising, Blume-Capel etc.). This compensation effect has not been investigated before in any discrete symmetric planar magnetic models. The present study explores the compensation behaviours in discrete symmetric planar magnetic model, i.e., clock model. We have chosen 6-state clock model (with anisotropy) in a trilayered structure. The ferromagnetic nearest neighbour intraplanar interaction and the antiferromagnetic interplanar nearest neighbour interactions are considered. The response of the system is investigated through the Monte Carlo methods using Metropolis single spin flip algorithm.
The sublattice magnetisations of three different layers, the total magnetisation, the susceptibility and the specific heat are calculated. These quantities are studied as functions of the temperature in equilibrium. For the isotropic case ($D=0$), the low temperature compensation (vanishing of the total magnetisation for nonzero values of sublattice magnetisations) and high temperature critical (each sublattice magnetisation vanish) behaviour are identified. These two special points (critical and compensation) are identified as two special values of the temperatures which generally provide the maxima of susceptibility and the specific heat.
We have thoroughly investigated the effects of anisotropy (single-site). How do the critical and compensation temperatures changes with the variation of anisotropy ? We have observed that both critical and compensation temperature increases with the increase of anisotropy and finally saturate. The comprehensive phase diagram is provided. The morphologies of the spin orientations (discrete) are investigated in four different regimes of the temperature.
The finite size analysis is done through the temperature dependence of fourth order Binder cumulant to determine the actual (in the thermodynamic limit) critical temperature. The susceptibility has been found to scale with the system size in a usual power law fashion. The exponent has been determined.
The q-state clock model maps to continuous symmetric XY model in the limit $q\to\infty$. It would be interesting to study the compensation behaviours, in the q-state clock model, by varying $q$. The extrapolation in the limit $q\to\infty$ should reach the compensation temperature for XY model. Moreover, the compensation in disordered (may be modeled by randomly distributed anisotropy or random field or both) system would really be an interesting and important investigation, as the continuous symmetric planar ferromagnet (XY model) showed interesting behaviours in the case of randomly distributed anisotropy[28] and random field[29].
Acknowledgements: IT acknowledges UGC JRF, Govt. of India, for financial support. We are thankful to the computational facilities provided by Presidency University, Kolkata. Data availability statement: Data will be available on reasonable request to Olivia Mallick.
Conflict of interest statement: We declare that this manuscript is free from any conflict of interest.
Funding statement: No funding was received, particularly to support this work.
Authors’ contributions: Olivia Mallick developed the code, collected the data, prepared the figures, analysed the results and wrote the manuscript. Muktish Acharyya conceptualised the problem, analysed the results and wrote the manuscript.